Thad Morgan, a motorcycle enthusiast, has been exploring the possibility of relaunching the Western Hombre brand of cycle that was popular in the 1930s. The retro-look cycle would be sold for $11,000 and at that price, Thad estimates 600 units would be sold each year. The variable cost to produce and sell the cycles would be $8,250 per unit. The annual fixed cost would be $1,320,000.

a. What is the break-even in unit sales?

b. What is the margin of safety in dollars?

c. What is the degree of operating leverage? (Round your answer to 2 decimal places.)

Thad is worried about the selling price. Rumors are circulating that other retro brands of cycles may be revived. If so, the selling price for the Western Hombre would have to be reduced to $9,500 to compete effectively. In that event, Thad would also reduce fixed expenses to $1,030,000 by reducing advertising expenses, but he still hopes to sell 600 units per year.

d. What would the net operating income be in this situation? (Negative amount should be indicated by a minus sign.)
Answer #1

Correct Answer:

Break Even point in unit sales

$ 480

Margin of safety in Dollars

$ 13,20,000

Degree of Operating Leverage

  5 Times

Working:

Working

Current

A

Sales price

$           11,000.00

B

Variable cost

$              8,250.00

C=A-B

Contribution margin

$              2,750.00

D

Units (Motorcycle)

600

E=C x D

Total contribution margin

$     16,50,000.00

F

Fixed costs

$     13,20,000.00

G=E-F

Net Income

$        3,30,000.00

I=F/C

Break Even point in unit sales

$                       480

J=(D -I) * A

Margin of safety ratio

$           13,20,000

K=E/G

Degree of Operating Leverage

$                      5.00

Requirement 2:

Net Income

$      (2,80,000.00)

Working:

Working

Proposed Automation

A

Sales price

$              9,500.00

B

Variable cost

$              8,250.00

C=A-B

Contribution margin

$              1,250.00

D

Units (Motorcycles)

600

E=C x D

Total contribution margin

$        7,50,000.00

F

Fixed costs

$      10,30,000.00

G=E-F

Net Income

$ - 2,80,000.00

End of answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks.